The following particulars relate to a symmetrical tangent cam operating a roller follower :-
Least radius = 36 mm, nose radius = 30 mm, roller radius = 21 mm, distance between cam
shaft and nose centre = 28 mm, angle of action of cam = 150
1
Expert's answer
2021-07-08T04:31:02-0400

We have,

Total lift + r1 = PO + r2

20 + 30 = d – 5 d

= 45 mm

"R = (r1^2- r2^2 =d^2-2r1.d.cos \alpha)/ 2(r1-r2-dcos\alpha"

"R= (30^2-5^2 =45^2-2*30*45*cos75)\/ 2(30-5-45cos75)"

R= 82.42mm

therefore PQ = R-r2=82.41-5 = 77.41mm

QO = R-ri = 82.41-30 = 52.41mm